OpenStudy (anonymous):

indicate the equation of the line through (2,-4) and having slope of 3/5

OpenStudy (anonymous):

_x__y=____

OpenStudy (anonymous):

whats the equation for a line?

OpenStudy (anonymous):

i just dont get that form they have it in....

OpenStudy (anonymous):

well I'm not exactly good at this, but couldn't you just plug it into y = mx + b

OpenStudy (anonymous):

so I think you might use this equation y - y1 = m(x - x1)

OpenStudy (anonymous):

plug the -4 into x1 and the 2 into y1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then plug the slope into m

OpenStudy (anonymous):

so y+4 = 3/5(x-2)

OpenStudy (anonymous):

then distribute the 3/5

OpenStudy (anonymous):

so y + 4 = 3/5(x) - 3/5(2)

OpenStudy (anonymous):

which is y + 4 = 3/5(x) - 6/5

OpenStudy (anonymous):

got that ..

OpenStudy (anonymous):

subtract 4 from both sides to isolate y

OpenStudy (anonymous):

so y = 3/5(x) - 6/5 - 4

OpenStudy (anonymous):

combine the terms and end up with: y = (3/5)x -26/5

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

hope it's right :O